OpenStudy (anonymous):

Convert 95° into radians.

jimthompson5910 (jim_thompson5910):

multiply that given angle in degrees by pi/180 to convert to radians

OpenStudy (anonymous):

thank you very much

jimthompson5910 (jim_thompson5910):

you're welcome, tell me what you get

OpenStudy (anonymous):

1.657

jimthompson5910 (jim_thompson5910):

keep it as a fraction though

jimthompson5910 (jim_thompson5910):

in terms of pi

jimthompson5910 (jim_thompson5910):

do you know what I mean?

OpenStudy (anonymous):

no I don't know how to do that:(

jimthompson5910 (jim_thompson5910):

ok I'll show you what I mean

jimthompson5910 (jim_thompson5910):

Multiply the given angle 95 degrees by pi/180 to get \[\large 95*\frac{\pi}{180}\] \[\large \frac{95\pi}{180}\] \[\large \frac{5*19\pi}{5*36} \ ... \ \text{ Factor}\] \[\large \frac{19\pi}{36} \ ... \ \text{ Cancel (to reduce)}\] So \[\large 95^{\circ} = \frac{19\pi}{36} \text{ radians}\]

OpenStudy (anonymous):

how did you get 19

OpenStudy (anonymous):

ohhh you factored

jimthompson5910 (jim_thompson5910):

yes 19*5 = 95

OpenStudy (anonymous):

\[180degree=\pi radians,1 degree=\frac{ \pi }{180 },95 degree=\frac{ 95\pi }{180 }=\frac{19\pi }{ 36 }radians\]